WAC Notes August 21 2019

From PREX Wiki
Jump to: navigation, search
Documentation
HOW TOs for shift crew
Expert Tools
All Expert Contacts



PREX Main << Weekly Analysis Coordinator << WAC Notes << Instructions

WAC Notes August 20 2019 << >> WAC Notes August 22 2019

August 21 2019

Daily Meeting BlueJeans

Production Run List - Slug 59 In (Wednesday and Thursday Owl)

  • Slug occured between Owl Shifts (Moller and Q^2 Studies In Between)
  • Run 4368 - Good - 4368 - RHRSMain detector moved by accident, s alvaged a few thousand events putting cut on event number, target degrading
  • Run 4369 - Good - 4369 - Target Degrading
  • Run 4370 - Good - 4370 - Target Degrading
  • Run 4372 - Good - 4372 - Move to Pb 5 Target, widths decreased
  • Run 4373 - Good - 4373 - Good
  • Run 4374 - Good - 4374 - Good
  • After Moller and Q^2
  • Run 4392 - Good - 4392 - Good
  • Run 4393 - Good - 4393 - Good (Compton tuning)
  • Run 4394 - Good - 4394 - Good (Compton tuning)
  • Run 4395 - Good - 4395 - Good
  • Run 4396 - Good - 4396 - Outlier in regressed us avg for first minirun - needs a cut. Beam steered during Compton tuning to reduce background rates
  • Run 4397 - Good - 4397 - Good
  • Run 4398 - Good - 4398 - Good
  • Run 4399 - Good - 4399 - Good. Some confusion during this r un. Beam moved which was reflected in epics and the data. The BPMs and the complements were seeing different things.
  • Run 4400 - Good - 4300 - Good
  • Run 4401 - Good - 4401 - Good. Same thing happened here as it did in Run 4399
  • Run 4402 - Good - 4402 - Good. Bad beam though.....
  • Run 4403 - Good - 4403 - Good.
  • Run 4404 - Good - 4404 - Good.
  • Run 4405 - Good - 4405 - Good
  • Run 4406 - Good - 4406 - Good. Run stopped CODA failure

Non-Production Runs and Bad Runs

  • Run 4371 - Junk, Empty Data
  • Run 4375 - Junk
  • Runs 4376-4382 Test BCM comparison to Moller test
  • Run 4383 - Test BCM comparison to Moller test changes to Pockel cell voltage
  • Runs 4384-4385 - Test PITA scan and PITA setting change
  • Run 4386 - Test BCM comparison
  • Runs 4387 - 4389 Test HWP scans
  • Run 4390 - Test RWHP sweep
  • Run 4391 - DAQ test
  • Run 4007 - Junk

Pedestal/Calibrations Updates

  • Channel map made for Run 4368 to cut out events where the detector moved.

Relevant Information

Other Status Updates